Company ABC offers a common stock that pays an annual dividend of $3 a share. The company has promised to maintain a constant dividend. How much are you willing to pay for one share of this stock if you want to earn a 10% return on your investment?


1
Expert's answer
2021-02-23T12:54:22-0500

X*0.1=3, then x=3/0.1=30